sarah is 7 years older than sheila. Find their present age if the product of sarah's age 6 years from now and sheila's age is 7 years ago is 69.

Answer by ikleyn(52788)

.

Let x be the Sarah age. Then the Sheila age is x-7 years.


Sarah's age 6 years from now is (x+6) years;

Sheila's age 7 years ago is/was ((x-7)-7) = x-14 years.


So, the number 69 is the product of factors  (x+6) and x-14 with the difference between them of 20.


There is only one such a pair in positive integer numbers: it is (23,3).


So,  x+6 = 23, which means that Sarah is 17 years old, while Sheila is 17-7 = 10 years old.    ANSWER

Solved.